Experiment is reduction of 9-fluorenone using sodium borohydride to give 9-fluorenol. I need theoretical yield and percent yield calculations

intial: Fluorenone mass: 0.206g

Final: Fluorenol mass: 0.169g
Answer #1

for given reaction

ОН

molecular mass of reactant = 180.21 g/mol

no of moles of reactant = 0.206 g/180.21 gmol-1 = 1.14 mmol

expected moles of product 1.14 mmol

expected mass of product = 1.14 mol x 182.22 gmol-1 x10-3 = 0.208 g

actual mass obtained = 0.169 g

yield = 0.169*100/0.208 = 81.13 %